which of the ordered pairs are solutions for the equation x-y=4?
(3,1), (0,-4), (-4,0), (-3-7) with instructions on what was done to get the answer

Answer by nidhi_mundra_5(2)

x-y=4----->put the value -4 for y and 0 for x
0-(-4)=4
0+4=4
4=4
hence the second pair is the correct option